OpenStudy (kmullis6):

WILL FAN AND MEDAL Solve for x: the quantity of (x + 16) over 3 = 3x x = -7 x = -13 x = 8 x = 2

OpenStudy (anonymous):

x = 2

OpenStudy (kmullis6):

@boggley can you explain this to me? i don't understand how you got the answer... /:

OpenStudy (anonymous):

are you still taking english?

OpenStudy (anonymous):

\[\frac{x+16}{3}= 3x \implies x+16 = (3)3x \implies x +16 = 9x \implies 16=8x.\] So \[\frac{16}{8}=\frac{8x}{8} \implies x =2.\]

OpenStudy (kmullis6):

@dummyguy ok thanks i get it now (: